# Theory, and Design Guidelines

### v1

@@ IV-A Convergence Guarantees of Deep Unfolded Optimizers
Parameter Coupling and Convergence Rates in LISTA and Its Variants Consider the sparse signal recovery problem arising from (4) and assume ${\boldsymbol{\Psi}}={\boldsymbol{I}}$ for simplicity. Let the observation be ${\boldsymbol{x}}={\boldsymbol{H}}{\boldsymbol{s}}_{\star}+{\boldsymbol{e}}$, where $\boldsymbol{s}$$_{\star}$ is the ground-truth sparse signal and $\boldsymbol{e}$} is observation noise. We estimate $\boldsymbol{s}$$_{\star}$ using the $k$-th layer output ${\boldsymbol{s}}^{(k)}$ of the LISTA network (13). Assuming that $\lVert{\boldsymbol{s}}_{\star}\rVert_{\infty}\leq B$, $\lVert{{\boldsymbol{s}}_{\star}}\rVert_{0}\leq\eta$, and $\lVert{{\boldsymbol{e}}}\rVert_{1}\leq\sigma$, Chen et al. [41] established the following: 1. If $\sigma=0$ and ${\boldsymbol{s}}^{(k)}\to{\boldsymbol{s}}_{\star}$ as $k\to\infty$, then under mild conditions the parameters must satisfy ${\boldsymbol{W}}^{k}_{2}-({\boldsymbol{I}}-{\boldsymbol{W}}^{k}_{1}{\boldsymbol{H}})\to 0,\qquad\beta_{k}\to 0.$ (18) 2. Building on (18), the authors proposed a coupled-parameter (CP) technique to reduce the number of learnable parameters by enforcing ${\boldsymbol{W}}^{k}_{2}\triangleq{\boldsymbol{I}}-{\boldsymbol{W}}^{k}_{1}{\boldsymbol{H}}$. They proved that, when $\eta$ is sufficiently small, there exists $\left\{{\boldsymbol{W}}^{k}_{1},\beta_{k}\right\}_{k=1}^{\infty}$ and constants $c,C>0$ that guarantees for all possible $({\boldsymbol{s}}_{\star},{\boldsymbol{e}})$: $\left\lVert{\boldsymbol{s}}^{(k)}-{\boldsymbol{s}}_{\star}\right\lVert_{2}\leq\eta B\exp(-ck)+C\sigma,\quad\forall k=1,2,\cdots.$ Furthermore, it was shown in [42] that for LISTA-CP, the weight matrices ${\boldsymbol{W}}^{k}_{1}$ need not be learned from data; it suffices to take them proportional to a matrix that minimizes the mutual coherence of $\boldsymbol{H}$}. [^1] The mutual coherence optimization problem is nonconvex, thus it is difficult to compute global minimizers in general. Under reasonable conditions, it is shown that this analytical choice (the resultant design is termed ALISTA) yields linear convergence, while the convergence rate of LISTA cannot be faster than linear, thus leading a near-optimality architecture with better parameter efficiency. Moreover, by introducing momentum term to ALISTA and instance-adaptive hyperparameter scheduling, the resulting unrolled optimizer can be guaranteed to achieve superlinear convergence under appropriate assumptions[43]. More recently, several studies have extended these ideas to broader sparse-recovery settings. For example, Yang et al. [44] analyze an unfolded architecture for multiple measurement vectors sparse recovery and derive CP-type relations and linear convergence guarantees akin to [41], together with analytic expressions for certain weights analogous to [42]. Yang et al. [45] unfold a proximal-gradient method for the Elastic Net, and establish CP-type relations and linear convergence.
@@ IV-B Generalization Error Bounds of Unfolded Networks
To assess performance on unseen data, we study the generalization gap $\mathcal{L}({\boldsymbol{\theta}})-\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})$. Since $\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})$ is random, we typically consider high-probability bounds of the generalization gap or bounds on the expected supremum $\mathbb{E}_{\mathcal{D}}\left[\sup_{{\boldsymbol{\theta}}}\left(\mathcal{L}({\boldsymbol{\theta}})-\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})\right)\right]$. Both types of bounds can be controlled by the Rademacher complexity defined as [50, Ch. 26]
@@
where $\Theta$ is the set of possible parameter values, $\left\{\epsilon_{t}\right\}_{t=1}^{\left\lvert\mathcal{D}\right\rvert}$ are i.i.d. Rademacher variables with $\mathbb{P}\left(\epsilon_{t}=1\right)=\mathbb{P}\left(\epsilon_{t}=-1\right)=\frac{1}{2}$. Hence, bounding the generalization gap reduces to upper-bounding $\mathfrak{R}_{\left\lvert\mathcal{D}\right\rvert}$. We collect several representative results for sparse-recovery–oriented unfolded networks, together with key proof ideas, in the box titled “Generalization Error Bounds of Unfolded Networks for Sparse Recovery” on Page IV-B. While such conclusions are stated for particular architectures, the analysis techniques often extend to many other unfolded designs.
@@ IV-C Optimization Guarantees for Unfolded Network Training
Similar results hold for standard feedforward DNNs; moreover, the authors demonstrates by theory and experiments that, under equal parameter budgets, unfolded architectures can interpolate (i.e., “memorize”) more samples than feedforward DNNs in the same setting. This suggests that unfolded structures are better matched to sparse recovery, rendering them a powerful architectural choice. Notably, the analysis techniques in this work may be applied to broader families of unfolded architectures in the overparameterized regime.
@@ V Comparative Study
Section III presented a set of design approaches and concrete methodologies for converting iterative optimizers into deep unfolded ML architectures. While all these methods originate from iterative optimization, they notably vary in their strengths, requirements, and implications. To highlight the interplay between the approaches and provide an understanding of how practitioners should prioritize one approach over the others, we next provide a comparative study. We divide this study into two parts: we first present a qualitative comparison, which pinpoints the conceptual differences between the approaches in light of the identified desired properties and challenges listed in Table I. We then detail a quantitative study for a representative scenario of iterative optimization based on the RPCA objective to capture the regimes in which each approach is expected to be preferable.
@@ V-A Qualitative Comparison
The different design methodologies of deep unfolding, reviewed in Section III, each provide distinct means of converting iterative optimization solvers into trainable ML architectures. While they share a common foundation in iterative optimization, their degree of abstraction, interpretability, and learning capacity vary notably. To provide a unified perspective, we next conduct a qualitative comparative study highlighting the relative strengths and trade-offs among the four considered unfolding approaches. The aspects considered in this comparison are derived from the desired properties and challenges of iterative optimization-based inference discussed in Section II. Specifically, the ability of unfolded optimizers to obtain the minimizer of their corresponding objective (i.e., P1) is discussed in Section IV, and both challenges C1 and C2, while originating from different reasons, indicate that the model should be robust to mismatched and approximated objectives. Accordingly, the aspects inspected are including: Adaptability (P2); Interpretability (P3); Scalability (P4); Abstractness (C1C2); Latency (C3); and Complexity (C4). The comparison detailed below is summarized in Table II.
@@
Learning objective parameters Requires (possibly simplified) retraining for adaptivity Fully interpretable Requires (possibly simplified) retraining to handle data with different dimensions Can support complex mappings that are based on the structure of the objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$ (though not on its parameters ${\boldsymbol{\theta}}^{\rm o}$) Supports rapid implementations via few fixed iterations Depending on the specific iteration operation, though complex operation can be relaxed with approximations whose distortion is mitigated via learning [19]
@@
End-to-end DNN Not adaptive Black-box Requires retraining and architecture change to handle data with different dimensions Can learn complex mappings and is invariant of $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$ Depending on number of layers and DNN compactness Typically involves complex DNNs
@@
The various deep unfolding methodologies differ substantially in their their ability to employ the same learned optimizer across multiple settings characterized by different objective parameters ${\boldsymbol{\theta}}^{\rm o}$. Methods based on learning hyperparameters and learning correction terms explicitly incorporate the objective parameters ${\boldsymbol{\theta}}^{\rm o}$ into their formulation, thereby inheriting the adaptability of classical iterative solvers. Once trained, they can typically be applied to new instances of the same optimization problem (e.g., different sensing matrices or regularization structures) without retraining. However, in the case of learned correction terms, the correctional DNN components may need fine-tuning when the operating conditions diverge substantially from those observed during training, as their internal representations are implicitly data-dependent. Methods that learn objective parameters or adopt a DNN inductive bias generally require some degree of retraining when the objective or data distribution changes. Still, the adaptation process is often more tractable than in black-box DNNs, since unfolded architectures preserve modularity and interpretability, that allowing selective retraining or even targeted adjustment of a subset of parameters [36], and support low-complexity continual learning [17, 35]. Among the two, learning objective parameters tend to require fewer trainable weights, making adaptation simpler and faster, while DNN inductive bias models offer greater flexibility at the cost of heavier retraining requirements.
@@
Among the examined methodologies, those that fully preserve the operation of the original iterative solver (namely, learning hyperparameters and learning objective parameters) offer complete interpretability. In these cases, every computation carried out within each iteration corresponds to a known step of the underlying optimization algorithm, and the intermediate features can be directly interpreted as gradual refinements of the estimated optimization variable. Moreover, when learning objective parameters, the learned values can be explicitly linked to the objective being optimized at each iteration, providing a transparent view of how data-driven adaptation affects the mathematical formulation of the problem. Methods that integrate neural components introduce some level of black-box behavior. When learning correction terms, interpretability is largely preserved since the architecture still follows the computational flow of the original iterative solver, and the learned DNN only contributes small corrective adjustments to selected intermediate quantities. This ensures that the high-level operation remains understandable, even if the precise effect of the correction is harder to analyze analytically. At the other end of the spectrum, the DNN inductive bias approach departs more significantly from the original solver, as each iteration is implemented through a neural module. Nonetheless, even in this highly abstract setting, partial interpretability is maintained: the exchanged features between iterations retain their meaning as progressively refined estimates of the decision variable, a property that is frequently exploited in the training and analysis of such architectures.
@@
As in P4, by scalability we refer to the ability of the learned model to generalize across problems of different dimensions, such as varying signal sizes or system configurations. Unfolded architectures that learn hyperparameters retain the full mathematical structure of the original iterative optimizer and therefore exhibit the strongest scalability among all approaches. Once trained, they can be applied to data of different dimensions without any retraining or structural modification, provided the underlying optimization model remains consistent. The remaining methodologies are typically more tied to the data dimensions used during training, as these dimensions often determine their parameterization and neural network structures. Nevertheless, their modular and interpretable nature allows more efficient adaptation than conventional black-box DNNs. In particular, dimensional adjustments can often be achieved through limited retraining or the use of modular hypernetworks that generate dimension-specific parameters for each component as in [37].
@@
Abstractness, in this context, refers to the ability of the unfolded optimizer to learn effective inference rules even when the original optimization objective does not accurately represent the true underlying task. Among the considered methodologies, learning hyperparameters offers only limited abstractness, as it retains the exact optimization structure and merely tunes the iterative dynamics. While the trainable nature of its hyperparameters introduces some flexibility compared to the conventional solver, its inference behavior remains tightly coupled to the original objective formulation. A similar property holds for learning correction terms, although the inclusion of neural correction modules introduces additional degrees of freedom. These corrections can adapt through supervised learning to partially compensate for modeling inaccuracies and mismatched objectives, providing a moderate level of abstraction while preserving interpretability. The remaining approaches provide increased abstractness by allowing the optimizer to deviate more freely from the original mathematical formulation. In learning objective parameters, each iteration can adapt its own learned objective, effectively redefining what is being optimized at each step to better align with the data-driven task. The DNN inductive bias approach achieves the highest degree of abstractness: by implementing each iteration as a DNN, the learned mappings are no longer constrained by the structure of the original solver, thus allowing the model to discover highly flexible and potentially more accurate inference rules directly from data.
@@
Latency refers to the time required to produce an inference result, which in the context of iterative optimization corresponds to the number of iterations or computational layers that must be executed sequentially. Unfolded architectures based on learning hyperparameters and learning objective parameters preserve the computational structure of the original solver but operate with a fixed and predefined number of iterations. As such, they can be designed to achieve low latency by setting the number of iterations to a small value while maintaining high performance through data-driven parameter tuning. Methods that incorporate neural components may introduce additional latency due to the need for a forward pass through the DNNs at each iteration. The degree of additional latency depends on the underlying hardware and implementation, as neural computations can often be highly parallelized and accelerated using dedicated inference engines. Despite this, these approaches remain inherently based on iterative solvers with a small and fixed number of iterations, and thus their overall latency typically remains low compared to traditional optimization methods that require numerous iterations to converge.
@@
Complexity captures the number of arithmetic operations required by the model during inference and directly depends on the computational structure of each iteration. Unfolded optimizers that faithfully inherit the operations of the original iterative solver, such as those based on learning hyperparameters or learning objective parameters, exhibit low complexity, as they operate with a limited and fixed number of iterations. Moreover, their trainable parameters can be optimized to compensate for approximations introduced when simplifying expensive operations, such as matrix inversions, thereby achieving favorable accuracy–complexity trade-offs [19]. As for the DNN-aided methodologies, i.e., learning correction terms and DNN inductive bias, their complexity is primarily determined by the size and structure of the neural networks employed within each iteration. Although the inclusion of DNNs increases the computational load relative to purely algorithmic solvers, the deep unfolding framework facilitates using compact, structured modules, whose number of parameters (and thus their computational complexity) remains significantly lower than that of end-to-end black-box DNNs.

### v2

@@ IV-A Convergence Guarantees of Deep Unfolded Optimizers
Parameter Coupling and Convergence Rates in LISTA and Its Variants Consider the sparse signal recovery problem arising from (4) and assume ${\boldsymbol{\Psi}}={\boldsymbol{I}}$ for simplicity. Let the observation be ${\boldsymbol{x}}={\boldsymbol{H}}{\boldsymbol{s}}_{\star}+{\boldsymbol{e}}$, where $\boldsymbol{s}$$_{\star}$ is the ground-truth sparse signal and $\boldsymbol{e}$} is observation noise. We estimate $\boldsymbol{s}$$_{\star}$ using the $k$-th layer output ${\boldsymbol{s}}^{(k)}$ of the LISTA network (13). Assuming that $\lVert{\boldsymbol{s}}_{\star}\rVert_{\infty}\leq B$, $\lVert{{\boldsymbol{s}}_{\star}}\rVert_{0}\leq\eta$, and $\lVert{{\boldsymbol{e}}}\rVert_{1}\leq\sigma$, Chen et al. [49] established the following: 1. If $\sigma=0$ and ${\boldsymbol{s}}^{(k)}\to{\boldsymbol{s}}_{\star}$ as $k\to\infty$, then under mild conditions the parameters must satisfy ${\boldsymbol{W}}_{k}^{2}-({\boldsymbol{I}}-{\boldsymbol{W}}_{k}^{1}{\boldsymbol{H}})\to 0,\qquad\beta_{k}\to 0.$ (18) 2. Building on (18), the authors proposed a coupled-parameter (CP) technique to reduce the number of learnable parameters by enforcing ${\boldsymbol{W}}_{k}^{2}\triangleq{\boldsymbol{I}}-{\boldsymbol{W}}_{k}^{1}{\boldsymbol{H}}$. They proved that, when $\eta$ is sufficiently small, there exists $\left\{{\boldsymbol{W}}_{k}^{1},\beta_{k}\right\}_{k=1}^{\infty}$ and constants $c,C>0$ that guarantee for all possible $({\boldsymbol{s}}_{\star},{\boldsymbol{e}})$: $\left\lVert{\boldsymbol{s}}^{(k)}-{\boldsymbol{s}}_{\star}\right\lVert_{2}\leq\eta B\exp(-ck)+C\sigma,\quad\forall k=1,2,\cdots.$ Furthermore, it was shown in [50] that for LISTA-CP, the weight matrices ${\boldsymbol{W}}_{k}^{1}$ need not be learned from data; it suffices to take them proportional to a matrix that minimizes the mutual coherence of $\boldsymbol{H}$}. [^1] The mutual coherence optimization problem is nonconvex, thus it is difficult to compute global minimizers in general. Under reasonable conditions, it is shown that this analytical choice (the resultant design is termed ALISTA) yields linear convergence, while the convergence rate of LISTA cannot be faster than linear, thus leading to a near-optimality architecture with better parameter efficiency. Moreover, by introducing a momentum term to ALISTA and instance-adaptive hyperparameter scheduling, the resulting unrolled optimizer can be guaranteed to achieve superlinear convergence under appropriate assumptions[51]. More recently, several studies have extended these ideas to broader sparse-recovery settings. For example, Yang et al. [52] analyze an unfolded architecture for multiple measurement vectors sparse recovery and derive CP-type relations and linear convergence guarantees akin to [49], together with analytic expressions for certain weights analogous to [50]. Yang et al. [53] unfold a proximal-gradient method for the Elastic Net, and establish CP-type relations and linear convergence.
@@ IV-B Generalization Error Bounds of Unfolded Networks
To assess performance on unseen data, we study the generalization gap $\mathcal{L}({\boldsymbol{\theta}})-\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})$. Since $\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})$ is random, we typically consider high-probability bounds of the generalization gap or bounds on the expected supremum $\mathbb{E}_{\mathcal{D}}\left[\sup_{{\boldsymbol{\theta}}}\left(\mathcal{L}({\boldsymbol{\theta}})-\mathcal{L}^{\rm ER}_{\mathcal{D}}({\boldsymbol{\theta}})\right)\right]$. Both types of bounds can be controlled by the Rademacher complexity defined as [58, Ch. 26]
@@
where $\Theta$ is the set of possible parameter values, $\left\{\epsilon_{t}\right\}_{t=1}^{\left\lvert\mathcal{D}\right\rvert}$ are i.i.d. Rademacher variables with $\mathbb{P}\left(\epsilon_{t}=1\right)=\mathbb{P}\left(\epsilon_{t}=-1\right)=\frac{1}{2}$. Hence, bounding the generalization gap reduces to bounding $\mathfrak{R}_{\left\lvert\mathcal{D}\right\rvert}$. We collect several representative results for sparse-recovery–oriented unfolded networks, together with key proof ideas, in the box titled “Generalization Error Bounds of Unfolded Networks for Sparse Recovery” on Page IV-B. While these are stated for particular architectures, the analysis techniques often extend to many other unfolded designs.
@@ IV-C Optimization Guarantees for Unfolded Network Training
Similar results hold for standard feedforward DNNs; moreover, the authors demonstrate by theory and experiments that, under equal parameter budgets, unfolded architectures can interpolate (i.e., “memorize”) more samples than feedforward DNNs in the same setting. This suggests that unfolded structures are better matched to sparse recovery, rendering them a powerful architectural choice. Notably, the analysis techniques in this work may be applied to broader families of unfolded architectures in the overparameterized regime.
@@ V Comparative Study
Section III presented a set of design approaches and concrete methodologies for converting iterative optimizers into deep unfolded ML architectures. Many deep unfolded architectures (particularly those incorporating neural modules within each iteration) can be viewed as end-to-end trained DNNs [65, 66] with a specific, algorithm-inspired architecture. Nonetheless, the deep unfolding paradigm itself is broader, being defined by the systematic transformation of an iterative optimization algorithm into a structured trainable model whose intermediate representations retain an explicit algorithmic interpretation.
The reviewed deep unfolding methods all originate from iterative optimization, still, they notably vary in their strengths, requirements, and implications. To highlight the interplay between the approaches and provide an understanding of how practitioners should prioritize one approach over the others, we next provide a comparative study. We divide this study into two parts: we first present a qualitative comparison, which pinpoints the conceptual differences between the approaches in light of the identified desired properties and challenges listed in Table I. We then detail a quantitative study for a representative scenario of iterative optimization based on the RPCA objective to capture the regimes in which each approach is expected to be preferable.
@@ V-A Qualitative Comparison
The different design methodologies reviewed in Section III, provide distinct means of converting iterative optimization solvers into trainable ML architectures. While they share a common foundation in iterative optimization, their degree of abstraction, interpretability, and learning capacity vary notably. To provide a unified perspective, we next conduct a qualitative comparative study highlighting the relative strengths and trade-offs. The aspects considered in this comparison are derived from the desired properties and challenges discussed in Section II. Specifically, the ability of unfolded optimizers to obtain the minimizer of their corresponding objective (i.e., P1) is discussed in Section IV, and both challenges C1 and C2, while originating from different reasons, indicate that the model should be robust to mismatched and approximated objectives. Accordingly, the aspects inspected are including: Adaptability (P2); Interpretability (P3); Scalability (P4); Abstractness (C1C2); Latency (C3); and Complexity (C4). The comparison detailed below is summarized in Table II.
@@
Learning objective parameters Requires (possibly simplified) retraining for adaptivity Partially interpretable, as one can identify the objective and solver hyperparameters associated with each iteration Requires (possibly simplified) retraining to handle data with different dimensions Can support complex mappings that are based on the structure of the objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$ (though not on its parameters ${\boldsymbol{\theta}}^{\rm o}$) Supports rapid implementations via few fixed iterations Depending on the specific iteration operation, though complex operation can be relaxed with approximations whose distortion is mitigated via learning [19]
@@
End-to-end DNN Not adaptive Black-box Some DNN architectures may also support variability in some dimensions, though in general, scalability requires retraining and architecture change Can learn complex mappings and is invariant of $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$ Depending on number of layers and DNN compactness Typically involves complex DNNs
@@
The various methodologies differ substantially in their ability to employ the same learned optimizer across multiple settings characterized by different objective parameters ${\boldsymbol{\theta}}^{\rm o}$. Methods based on learning hyperparameters and learning correction terms explicitly incorporate ${\boldsymbol{\theta}}^{\rm o}$ into their formulation, thereby inheriting the adaptability of classical iterative solvers. Once trained, they can typically be applied to new instances of the same optimization problem. However, in the case of learned correction terms, the correctional DNN components may need fine-tuning when the operating conditions diverge substantially from those observed during training, as their internal representations are implicitly data-dependent. Methods that learn objective parameters or adopt a DNN inductive bias generally require some degree of retraining when the objective or data distribution changes. Still, the adaptation process is often more tractable than in black-box DNNs, since unfolded architectures preserve modularity and interpretability, which allows selective retraining or even targeted adjustment of a subset of parameters [44], and support low-complexity continual learning [17, 43]. Among the two, learning objective parameters tend to require fewer trainable weights, making adaptation simpler and faster.
@@
Among the examined methodologies, those that fully preserve the operation of the original iterative solver (namely, learning hyperparameters) offer complete interpretability: Every computation carried out within each iteration corresponds to a known step of the underlying optimization algorithm, and the intermediate features are gradual refinements of the estimated optimization variable. Moreover, when learning objective parameters, the learned values can be linked to the objective being optimized at each iteration, providing some level of interpretability as to how data-driven adaptation affects the mathematical formulation of the problem, though there is not necessarily a formal relationship between the layer-wise objectives and the original formulation. Methods that integrate neural components introduce some level of black-box behavior. When learning correction terms, interpretability is largely preserved since the architecture still follows the computational flow of the original iterative solver. DNN inductive bias departs more significantly from the original solver, as each iteration is implemented through a neural module. Nonetheless, even in this highly abstract setting, partial interpretability is maintained: the exchanged features between iterations retain their meaning as progressively refined estimates of the decision variable, a property that can be exploited in training and analysis.
@@
As in P4, by scalability we refer to the ability of the learned model to generalize across problems of different dimensions, such as varying signal sizes or system configurations. Specific end-to-end DNNs may also support dimension variability, though this depends on architectural design, and often the architecture depends on the data dimensions. Unfolded architectures that learn hyperparameters retain the full structure of the original iterative optimizer and therefore exhibit strong scalability. Once trained, they can be applied to data of different dimensions without any retraining or structural modification, provided the underlying optimization model remains consistent. The remaining methodologies are typically more tied to the data dimensions used during training, as these dimensions often determine their parameterization. Nevertheless, their modular and interpretable nature allows more efficient adaptation than black-box DNNs. In particular, dimensional adjustments can often be achieved through limited retraining or the use of modular hypernetworks that generate dimension-specific parameters for each component as in [45].
@@
Abstractness, in this context, refers to the ability to learn effective inference rules even when the original optimization objective does not accurately represent the true underlying task. Among the considered methodologies, learning hyperparameters offers only limited abstractness, as it retains the exact optimization structure. While the trainable nature of its hyperparameters introduces some flexibility compared to the conventional solver, its inference behavior remains tightly coupled to the original objective formulation. A similar property holds for learning correction terms, although the inclusion of neural correction modules introduces additional degrees of freedom. The remaining approaches provide increased abstractness by allowing the optimizer to deviate more freely from the original mathematical formulation. In learning objective parameters, each iteration can effectively redefine what is being optimized at each step to better align with the data-driven task. DNN inductive bias achieves the highest degree of abstractness: by implementing each iteration as a DNN, the learned mappings are no longer constrained by the structure of the original solver, thus allowing the model to discover highly flexible inference rules directly from data.
@@
In the context of iterative optimization, latency corresponds to the number of iterations or computational layers executed sequentially. Unfolded architectures based on learning hyperparameters and learning objective parameters preserve the computational structure of the original solver but operate with a fixed and predefined number of iterations. As such, they can achieve low latency by setting the number of iterations to a small value while maintaining high performance through parameter learning. Methods that incorporate neural components may introduce additional latency due to the DNN forward pass at each iteration. The degree of additional latency depends on the underlying hardware and implementation, as neural computations can often be accelerated using dedicated inference engines. Despite this, these approaches remain inherently based on iterative solvers with a small and fixed number of iterations, and thus their overall latency typically remains low compared to traditional optimization methods that require numerous iterations to converge.
@@
Complexity captures the number of arithmetic operations required by the model during inference. Unfolded optimizers that faithfully inherit the operations of the original iterative solver, such as those based on learning hyperparameters or learning objective parameters, exhibit low complexity, as they operate with a limited and fixed number of iterations. Moreover, their trainable parameters can be optimized to compensate for approximations introduced when simplifying expensive operations, such as matrix inversions, thereby achieving favorable accuracy–complexity trade-offs [19]. As for the DNN-aided methodologies, i.e., learning correction terms and DNN inductive bias, their complexity is primarily determined by the size and structure of the neural networks employed within each iteration. Still, the deep unfolding framework facilitates using compact, structured modules, whose number of parameters (and thus their computational complexity) remains significantly lower than that of end-to-end black-box DNNs.